# A Mixed-Methods Study on the Implications of Unsafe Rust for Interoperation, Encapsulation, and Tooling

### v3

@@ 8. Conclusion
The Rust programming language provides static safety guarantees. However, Rust is frequently used to interoperate with other languages. To call foreign functions, developers need to use Rust’s unsafe features, which bypass its safety restrictions. These features can introduce bugs and security vulnerabilities if they are used incorrectly. Few Rust-specific development tools provide comprehensive support for unsafe code, and even fewer support multi-language applications. Developers will need new solutions for testing, bug finding, and verification in this context. We conducted a mixed-methods, exploratory study to determine which interventions will be effective for developers who use unsafe code. We interviewed 19 developers who regularly wrote or edited unsafe code and created a survey that reached an additional 160 developers who engaged with unsafe code in some capacity. We examined the challenges associated with interoperation, limitations in tooling, developers’ motivations for using unsafe code, and how they reason about encapsulation.
The majority of participants used unsafe code to call foreign functions. Certain foreign libraries had aliasing and concurrency patterns that conflicted with the expectations of Rust’s type system, making them difficult to encapsulate. However, participants did not have a reliable solution for determining if their encapsulations were sound. Miri, a Rust interpreter, was the de facto solution for detecting undefined behavior, but participants who had used it before were deterred from using again it due to its slow performance and lack of support for key features. Most participants were motivated to use unsafe code by necessity, but they also cited performance and ergonomics as motivations. The majority followed best practices such as encapsulating and documenting unsafe code, but many were somewhat uncertain about their design patterns, and few felt that Rust’s documentation was always adequate for understanding how to use unsafe code correctly.
To resolve this uncertainty, our participants will need verification and analysis tools that are capable of providing guarantees of soundness and have robust support for multi-language applications. We expect that tools which integrate with Miri or use preexisting documentation or can generate specifications from documentation will be uniquely effective at reaching the majority of our audience. Any interventions should consider that developers may be motivated by a combination of necessity, performance, and ergonomics.

### v4

@@ 8. Conclusion
The Rust programming language provides static safety guarantees. However, Rust is frequently used to interoperate with other languages. To call foreign functions, developers need to use Rust’s unsafe features, which bypass its safety restrictions. These features can introduce bugs and security vulnerabilities if they are used incorrectly. Few Rust-specific development tools provide comprehensive support for unsafe code, and even fewer support multilanguage applications. Developers will need new solutions for testing, bug finding, and verification in this context. We conducted a mixed methods, exploratory study to determine which interventions will be effective for developers who use unsafe code. We interviewed 19 developers who regularly wrote or edited unsafe code and created a survey that reached an additional 160 developers who engaged with unsafe code in some capacity. We examined the challenges associated with interoperation, limitations in tooling, developers’ motivations for using unsafe code, and how they reason about encapsulation.
The majority of participants used unsafe code to call foreign functions. Certain foreign libraries had aliasing and concurrency patterns that conflicted with the expectations of Rust’s type system, making them difficult to encapsulate. However, participants did not have a reliable method to determine if their encapsulations were sound. Miri, a Rust interpreter, was the most popular solution for detecting undefined behavior, but participants who had used it before were deterred from using it again because of its slow performance and lack of support for key features. Most participants were motivated to use unsafe code by necessity, but they also cited performance and ergonomics as motivations. The majority followed best practices, such as encapsulating and documenting unsafe code, but many were somewhat uncertain about their design patterns, and few felt that Rust’s documentation was always adequate for understanding how to use unsafe code correctly.
Our participants would benefit from new static and dynamic analysis tools to assist with interoperation and with finding Rust-specific forms of undefined behavior. Dynamic analysis tools could improve on Miri to provide better performance, support for inline assembly, and the capability to detect aliasing violations triggered by operations that span foreign function boundaries. The Rust community should also continue to invest in improving and expanding documentation on unsafe code, so that longstanding issue threads can become integrated into standard community resources. A useful direction for future work would be to synthesize the methods used by Astrauskas et al. (Astrauskas et al., 2020) and Evans et al. (Evans et al., 2020), which provide different, complementary perspectives on the distribution of unsafe code. Future studies should consider the use of unsafe code in unpublished but publicly available crates, as well as the extent to which Rust libraries are exposed to unsafe operations through foreign function calls. The results of such a study would provide an excellent opportunity to triangulate the results described in this paper.
Acknowledgements.
We thank the Rust Community for their engagement and support, as well as Jenny Liang and Courtney Miller for their feedback on our study design.